OpenStudy (anonymous):

five children are arranged in a row of swings. how many different arrangements are possible?

OpenStudy (anonymous):

5!=120. The first child can choose all 5 swings, the 2nd child 4 swings, 3rd childe 3 swings, etc. so it is 5*4*3*2*1=120.

OpenStudy (anonymous):

@rrrrr i am confused about the 5! ? that doesn't make sense. how do i sovle that?

jimthompson5910 (jim_thompson5910):

5! just means you start with 5 and multiply it with 4, 3, 2, 1 to get 120

jimthompson5910 (jim_thompson5910):

another example 8! = 8*7*6*5*4*3*2*1 in general, you start with the number and you move down until you hit 1 (multiplying each number)

OpenStudy (anonymous):

5! is five factorial which basically means that it is 5*4*3*2*1. 4! is four factorial which means that it is 4*3*2*1. and such and such

OpenStudy (anonymous):

and how do u solve 9! ??

jimthompson5910 (jim_thompson5910):

8! = 8*7*6*5*4*3*2*1 9! = 9*8*7*6*5*4*3*2*1 etc etc

OpenStudy (anonymous):

9! is 9*8*7*6*5*4*3*2*1. n! is n*n-1*n-2*n-3*n-4*n-5...*5*4*3*2*1.

OpenStudy (anonymous):

okay then. thankyou

OpenStudy (anonymous):

youre welcome